If the distance from the earth to the sun is 9.3x10⁷ miles and the distance from Neptune to the sun is 2.8x10⁹ miles, the distance from the earth to the sun is what percent of the distance from Neptune to the sun?
A) 2.7%
B) 2.8%
C) 2.9%
D) 3.1%
E) 3.3%

Answer by math_tutor2020(3817)


I'll write something like 10⁷ as 10^7

A = distance from earth to sun
A = 9.3 * 10^7 miles approximately
A = 93 million miles approximately

B = distance from Neptune to sun
B = 2.8 * 10^9 miles approximately
B = 2.8 billion miles approximately

Divide the two values (small/large) to get
A/B = (9.3 * 10^7)/(2.8 * 10^9)
A/B = (9.3/2.8) * (10^7)/(10^9)
A/B = 3.321 * 10^(7-9)
A/B = 3.321 * 10^(-2)
A/B = 0.03321
A/B = 3.321%
A/B = 3.3%
The values are approximate.

Answer: E) 3.3%

Side note: The distances are approximate since the sun and planets are moving all the time, and we don't have circular orbits (instead they are elliptical ones).
The distances represent averages.
Also, even if everything held perfectly still, then we'd still have approximations considering the vast magnitude of each number.